# CyberRAG: An agentic RAG cyber attack classification and reporting tool

### v1

@@ Abstract
Intrusion Detection and Prevention Systems (IDS/IPS) in large enterprises can generate hundreds of thousands of alerts per hour, overwhelming security analysts with logs that demand deep, rapidly evolving domain expertise. Conventional machine-learning detectors trim the alert volume but still yield high false-positive rates, while standard single-pass Retrieval-Augmented Generation (RAG) pipelines often retrieve irrelevant context and fail to justify their predictions. To overcome these shortcomings, we present CyberRAG, a modular, agent-based RAG framework that delivers real-time classification, explanation, and structured reporting for cyber-attacks. A central LLM agent orchestrates (i) a pool of fine-tuned specialized classifiers, each tailored to a distinct attack family; (ii) tool adapters for enrichment and alerting; and (iii) an iterative retrieval-and-reason loop that continuously queries a domain-specific knowledge base until the evidence is both relevant and self-consistent. Unlike traditional RAG systems, CyberRAG embraces an agentic design that enables dynamic control flow and adaptive reasoning. This agent-centric architecture refines its threat labels and natural-language justifications autonomously, reducing false positives and enhancing interpretability. The framework is fully extensible: new attack types can be supported by simply adding a classifier without retraining the core agent. CyberRAG has been evaluated on SQL Injection, XSS, and SSTI benchmarks, achieving over 94% accuracy per class and pushing final classification accuracy to 94.92% through semantic orchestration. Generated explanations score up to 0.94 in BERTScore and 4.9/5 in GPT-4-based expert evaluation, with robustness preserved against adversarial and unseen payloads. These results show that agentic, specialist-oriented RAG can pair high detection accuracy with trustworthy, SOC-ready prose, offering a practical and scalable path toward semi-autonomous cyber-defence workflows.
@@ 1 Introduction
The cybersecurity landscape has advanced considerably, moving from manual, expert-driven processes to increasingly automated and intelligent systems. However, the interpretation and response to cyber threats remains largely semi-automated and dependent on human expertise, particularly in large-scale enterprise environments. Intrusion Detection and Prevention Systems (IDS/IPS) continue to be foundational to network defense [1], but they generate massive volumes of alerts, often hundreds of thousands per hour, many of which require expert validation. Although machine learning-based detectors can reduce this burden, they often suffer from high false positive rates and limited interpretability [2].
In recent years, large language models (LLMs) have gained traction in cybersecurity due to their ability to interpret threat data and support analysts through natural language reasoning. A key advancement in this area is RAG, which enriches model input with relevant context retrieved from external data sources. This combination improves the quality and grounding of responses. Still, most RAG implementations retrieve context only once before generating output, lacking the ability to refine queries, reason iteratively, or dynamically adapt to complex situations.
@@
Designed to integrate seamlessly with existing IDS infrastructures, CyberRAG activates upon receiving a flagged alert and autonomously processes the associated network traffic. It identifies the likely attack category, retrieves supporting contextual information, and generates a comprehensive report in natural language that describes the threat and suggests mitigation steps. An integrated LLM-powered assistant enables interactive querying of the report, providing analysts with deeper insights or recommended remediation actions. The system is easily customizable: organizations can expand its internal knowledge base using internal documents, architectural diagrams, or policy files—without the need to fine-tune the agent or classifiers.
In a user study, human analysts rated CyberRAG’s reports as clear, well-structured, and actionable in 93% of cases. Moreover, in a pilot deployment on live enterprise traffic, CyberRAG reduced triage time by approximately 45%, significantly accelerating the response process. These results demonstrate that an agent-based, specialist-driven RAG framework can deliver both high detection performance and interpretable, SOC-ready outputs, offering a practical step toward semi-autonomous cybersecurity operations.
The remainder of this paper is organized as follows: Section 2 introduces the topic by presenting foundational cybersecurity concepts, such as web-based attacks, LLMs, Ai agents and agentic Rag. Section 7 discusses the related work. Our methodology is detailed in Section 3, followed by a discussion presenting the Knowledge Bases and the datasets in Section 4. The results are shown in Section 5 and discussed in Section 6. Finally, conclusions are drawn in Section 8.
@@ 2 Background
The tool we propose is a custom-built solution designed to assist security operators in identifying and resolving vulnerabilities that commonly arise during normal system operations. It leverages well-known web-based attack vectors to simulate real-world threats and incorporates a collection of established concepts such as Large Language Models (LLMs), Intrusion Detection Systems, Agents and agentic RAG which serve as foundational pillars for our approach.
These technologies are central to the objectives we aim to achieve and directly influence the design and implementation of our tool. In the following section, we provide an overview of the key technologies and methodologies referenced throughout this work, setting the stage for the detailed presentation of our system.
@@ 2.2 Cyber-attacks
Web-based attacks are among the most dangerous forms of cyber-attacks, as they allow unprivileged users, such as casual visitors to a website, to gain an initial foothold within a corporate network. Once inside, attackers can exploit this access to perform lateral movements across systems, gradually escalating their privileges. By chaining multiple stages of lateral movement and privilege escalation, an attacker can eventually gain full control over the entire system. The ultimate objective of these attacks is either to seize control of the target machine—typically through code injection techniques like SSTI—or to exfiltrate sensitive data, as seen in SQL injection and Cross-Site Scripting (XSS) attacks. These representative attack vectors have been selected to evaluate the capabilities of our tool and are discussed in detail in the remainder of this section.
@@ 2.2.2 Server-Side Template Injection (SSTI)
Many modern web applications utilize template engines to dynamically generate HTML by embedding user inputs into predefined templates. If the application fails to properly sanitize or validate these inputs, the template engine may interpret and execute them as code. This can lead to serious security consequences, including unauthorized data access, exfiltration of sensitive information, or even full remote code execution (RCE) on the host server.
@@ 2.2.3 SQL Injection
An SQL injection is an attack that occurs when an attacker is able to insert and execute malicious code within a vulnerable SQL query, as discussed in [10]. These attacks are commonly used to bypass login mechanisms or to exfiltrate sensitive information from databases.
@@ 2.3 Large Language Models
Large Language Models (LLMs) constitute a transformative advancement within the broader landscape of artificial intelligence, with particular salience in the subfield of Natural Language Processing (NLP). Their capacity for generalized learning and contextual adaptability has engendered substantial paradigm shifts across a range of domains, including but not limited to cybersecurity, biomedical informatics, and legal analytics. These models have markedly enhanced the efficiency with which extensive corpora of unstructured textual data are processed and interpreted. Furthermore, LLMs have played a pivotal role in the evolution of generative artificial intelligence, enabling the automated synthesis of semantically coherent and contextually appropriate content across diverse application areas [11, 12, 13].
At the architectural level, LLMs are predominantly grounded in Transformer-based frameworks, which operationalize self-attention mechanisms to capture intricate linguistic dependencies and long-range contextual relationships. The training paradigm typically involves a two-phase process: an initial pre-training stage on large-scale, unlabeled datasets, followed by task-specific fine-tuning on curated, domain-relevant corpora [14, 15].
Large Language Models can be broadly divided into two principal architectural families: encoder-based models and decoder-based models, each optimized for different downstream objectives. Encoder-based models, such as BERT and RoBERTa, are tailored for discriminative tasks requiring bidirectional understanding, making them ideal for classification, semantic matching, and anomaly detection. In this work, such models are leveraged as specialized agents, each focusing on the detection of a particular type of cyber threat. Conversely, decoder-based models—exemplified by the GPT and LLaMA families—are autoregressive and optimized for generative tasks, such as explanation synthesis and narrative generation. These models are integrated into the Retrieval-Augmented Generation (RAG) pipeline, where they generate context-aware threat descriptions informed by both classifier outputs and external knowledge. This dual-model strategy capitalizes on the complementary strengths of both families to achieve accuracy, explainability, and robustness in threat classification and reasoning.
A growing body of literature has examined the deployment of LLMs within the cybersecurity domain. Notably, the integration of Retrieval-Augmented Generation (RAG) techniques has been shown to enhance both the factual accuracy and contextual relevance of model output. These hybrid approaches leverage external knowledge repositories, such as structured databases or document collections, to supplement the model’s internal representations and mitigate hallucination phenomena [16, 17].
Despite these advancements, general-purpose LLMs frequently exhibit limitations when confronted with highly specialized or adversarial contexts, such as the identification and characterization of specific cyber threats (e.g., SQL Injection, Cross-Site Scripting). These deficiencies underscore the imperative for more modular domain-adapted architectures capable of capturing the nuanced semantics and operational signatures inherent to such attack vectors [18, 19].
To operationalize this need for specialization, consider a sequence of $n$ input tokens representing potentially malicious content. Each token is first mapped to a real-valued vector of dimension $d$ via a learnable embedding function, thereby projecting discrete inputs into a continuous vector space that encodes both syntactic and semantic properties. Positional encodings are then added to preserve token order. These enriched representations pass through successive self-attention layers, which enable the model to capture long-range dependencies and subtle contextual cues, even when the inputs are obfuscated or adversarial in nature. The resulting contextualized embeddings support downstream tasks such as high-precision classification, semantic similarity estimation, and domain-specific language understanding, providing the foundation for more robust and interpretable cybersecurity applications.
@@ 2.4 AI Agents
The term autonomy refers to the capability and requisites necessary to determine the appropriate course of action to achieve a specified objective. An intelligent agent (AI Agent) is characterized by its ability to perceive its environment, respond to changes within it, initiate actions, and engage with other systems, which may include other agents or human users. One of the core principles foundational to AI agents is the concept of memory. Effective memory management improves an agent’s capability in maintaining contextual awareness. It enables the agent to draw on previous experiences effectively, thereby facilitating the development of incrementally informed decision-making abilities as time progresses.
As noted in [21], the emergence of LLMs represents another moment of progress in the realization of AI agents. Substantial advances in this direction have culminated in the emergence of LLM agents. In particular, LLM agents use LLM as reasoning and planning cores to decide the control flow of an application, while maintaining the characteristics of traditional AI agents. LLM agents enhanced the LLMs by allowing them to invoke external tools to solve specific tasks, such as mathematical calculations or code execution. In the end, the LLM can decide whether the generated answer is sufficient or if more work is needed.
@@ 2.5 Agentic RAG
Although general-purpose large language models (LLMs) and their associated agents offer remarkable versatility, they frequently lack the in-depth domain-specific expertise essential for addressing intricate and specialized problems. One potential solution to this limitation involves retraining or fine-tuning the model; however, undertaking such procedures can be both financially demanding and resource consuming. An alternative strategy to circumvent these challenges emerges from the RAG paradigm. This strategy enhances LLMs by connecting them to external domain-specific knowledge sources, such as document collections, through a retriever component. Functioning like a search engine, the retriever extracts relevant information from this external repository and integrates it directly into the LLM’s context. This provides the model with the necessary data to effectively answer complex requests.
@@ 3.1 Classification Tool
PD001 1’ and 3580 = ( select count ( * ) from domain.domains as t1, domain.columns as t2, domain.tables as t3 ) --- - 0.9999 0.3956 0.0673
PD002 1’’}}{{1016814002+3472965455}} {{’bo’}} {#comment#}{raw ’bo’.join(’7n’) %} {{’7n’}}{{3140320242+4078937248} 0.3999 0.9997 0.3830
@@ 3.2 RAG Tool
The core LLM engine leverages the semantic representation to automatically generate a natural language query that reflects the identified attack type, salient payload characteristics, and relevant contextual indicators. This query is submitted to a semantic search engine that indexes curated cybersecurity resources, including: (i) vulnerability databases (e.g., CVE, NVD), (ii) technical documentation (e.g., OWASP, MITRE ATT&CK), (iii) scientific literature and incident reports.
@@ 3.4 User Chat
Although typically hidden, some LLM engines provide access to their internal reasoning processes. In this example, DeepSeek-r1:8b is used, and its reasoning is presented below.

### v2

@@ Abstract
Intrusion Detection and Prevention Systems (IDS/IPS) in large enterprises can generate hundreds of thousands of alerts per hour, overwhelming analysts with logs requiring rapidly evolving expertise. Conventional machine-learning detectors reduce alert volume but still yield many false positives, while standard Retrieval-Augmented Generation (RAG) pipelines often retrieve irrelevant context and fail to justify predictions. We present CyberRAG, a modular agent-based RAG framework that delivers real-time classification, explanation, and structured reporting for cyber-attacks. A central LLM agent orchestrates: (i) fine-tuned classifiers specialized by attack family; (ii) tool adapters for enrichment and alerting; and (iii) an iterative retrieval-and-reason loop that queries a domain-specific knowledge base until evidence is relevant and self-consistent. Unlike traditional RAG, CyberRAG adopts an agentic design that enables dynamic control flow and adaptive reasoning. This architecture autonomously refines threat labels and natural-language justifications, reducing false positives and enhancing interpretability. It is also extensible: new attack types can be supported by adding classifiers without retraining the core agent. CyberRAG was evaluated on SQL Injection, XSS, and SSTI, achieving over 94% accuracy per class and a final classification accuracy of 94.92% through semantic orchestration. Generated explanations reached 0.94 in BERTScore and 4.9/5 in GPT-4-based expert evaluation, with robustness preserved against adversarial and unseen payloads. These results show that agentic, specialist-oriented RAG can combine high detection accuracy with trustworthy, SOC-ready prose, offering a flexible path toward partially automated cyber-defense workflows.
@@ 1 Introduction
The cybersecurity landscape has advanced considerably, moving from manual expert-driven processes to increasingly automated and intelligent systems [1]. However, the interpretation and response to cyber threats remains largely semi-automated and dependent on human expertise, particularly in large-scale enterprise environments. Intrusion Detection and Prevention Systems (IDS/IPS) continue to be foundational to network defense [2], but they generate massive volumes of alerts, often hundreds of thousands per hour, many of which require expert validation. Although machine learning-based detectors can reduce this burden, they often suffer from high false positive rates and limited interpretability [3].
Moreover, the output produced by current IDS/IPS is typically presented as raw log strings: highly informative, yet difficult to read and interpret. Logs are static artifacts that lack interactivity and do not provide the possibility for clarification or contextualization. As a result, when an alert is triggered, highly skilled analysts must manually investigate the underlying event, potentially diverting the attention of security teams from truly critical threats.
A promising direction to address the aforementioned challenges lies in leveraging LLMs to support the analysis of logs and alerts, enhancing both the readability of results and the analyst’s ability to make informed decisions. In recent years, large language models (LLMs) have gained traction in cybersecurity due to their ability to interpret threat data and support analysts through natural language reasoning [4]. A key advancement in this area is RAG [5], which enriches model input with relevant context retrieved from external data sources. This combination improves the quality and grounding of responses [6]. Still, most RAG implementations retrieve context only once before generating output, lacking the ability to refine queries, reason iteratively, or dynamically adapt to complex situations.
@@
Designed to integrate seamlessly with existing IDS infrastructures, CyberRAG activates upon receiving a flagged alert and autonomously processes the associated network traffic. It identifies the likely attack category, retrieves supporting contextual information, and generates a comprehensive report in natural language that describes the threat and suggests mitigation steps. An integrated LLM-powered assistant enables interactive querying of the report, providing analysts with deeper insights or recommended remediation actions. The system is easily customizable: organizations can expand its internal knowledge base using internal documents, architectural diagrams, or policy files, without the need to fine-tune the agent or classifiers.
The remainder of this paper is organized as follows: Section 2 introduces the topic by presenting foundational cybersecurity concepts, such as web-based attacks, and important LLMs notions including AI agents and Agentic RAG. Section 2.6 discusses the related work. Our methodology is detailed in Section 3, followed by a discussion presenting the Knowledge Bases and the datasets in Section 4. The results are shown in Section 5 and discussed in Section 6. Finally, conclusions are drawn in Section 7.
@@ 2 Background
The growing interconnection of web services has made the detection and mitigation of sophisticated cyber-attacks increasingly critical. Organizations today face threats ranging from credential theft to large-scale offensives, such as Distributed Denial of Service (DDoS) [9] attacks and Advanced Persistent Threats (APTs). Among the most critical vulnerabilities, web applications represent a prime target for malicious actors who exploit browser flaws, hijack user sessions, or inject harmful code.
CyberRAG leverages well-known web-based attack vectors to simulate real-world threats and incorporates a collection of established concepts such as LLMs, IDS, AI Agents and Agentic RAG technology which serve as foundational pillars for our approach. These technologies are central to the objectives we aim to achieve and directly influence the design and implementation of our tool. In the following, we provide an overview of the key technologies and methodologies referenced throughout this work.
@@ 2.2 Cyber-attacks
Web-based attacks are among the most dangerous forms of cyber-attacks, as they allow unprivileged users, such as casual visitors to a website, to gain an initial foothold within a corporate network. Once inside, attackers can exploit this access to perform lateral movements across systems, gradually escalating their privileges. By chaining multiple stages of lateral movement and privilege escalation, an attacker can eventually gain full control over the entire system. The ultimate objective of these attacks is either to seize control of the target machine, typically through code injection techniques like SSTI, or to exfiltrate sensitive data, as seen in SQL injection and Cross-Site Scripting (XSS) attacks. These representative attack vectors have been selected to evaluate the capabilities of our tool and are discussed in detail in the remainder of this section.
@@ 2.2.2 Server-Side Template Injection (SSTI)
Many modern web applications utilize template engines to dynamically generate HTML by embedding user inputs into predefined templates. If the application fails to properly sanitize or validate these inputs, the template engine may interpret and execute them as code. This can lead to serious security consequences, including unauthorized data access, exfiltration of sensitive information, or even full remote code execution (RCE) on the server.
@@ 2.2.3 SQL Injection
A SQL injection is an attack that occurs when an attacker is able to insert and execute malicious code within a vulnerable SQL query, as discussed in [15]. These attacks are commonly used to bypass login mechanisms or to exfiltrate sensitive information from databases.
@@ 2.3 Large Language Models
Large Language Models have become one of the most significant advances in Natural Language Processing (NLP), with growing applications in areas such as cybersecurity, medicine, and law [16, 17, 18, 19, 20]. These models are typically based on the Transformer architecture [19], which makes it possible to capture long-range dependencies through self-attention mechanisms. Their training usually follows a two-phase paradigm: a large-scale pre-training stage on general-purpose corpora, followed by task-specific fine-tuning on curated datasets. This combination enables LLMs to learn general language representations while remaining adaptable to specialized domains.
From an architectural perspective, LLMs are commonly distinguished into two broad families: encoder-based models and decoder-based models. While they share the same Transformer foundation, their objectives and downstream applications differ substantially. Taken together, the two families highlight complementary strengths: encoder models excel at precise and context-aware classification, while decoder models provide detailed and interpretable explanations.
2.3.1 Encoder-based Models
Encoder-based models, exemplified by BERT [17] and RoBERTa [20], are primarily designed for discriminative tasks that require a bidirectional understanding of text. By processing input in both directions, they capture fine-grained semantic relations, which makes them particularly effective for classification, semantic matching, and anomaly detection. In our system, such models are employed as specialized classifiers, each focusing on a distinct type of cyber-attack (e.g., SQL Injection, XSS, SSTI). This specialization allows the classifiers to capture the unique linguistic patterns and syntactic features associated with each attack vector.
2.3.2 Decoder-based Models
Decoder-based models, such as GPT [16] and the LLaMA family [21], follow an autoregressive objective, predicting the next token in a sequence. This makes them naturally suited for generative tasks such as text completion, explanation synthesis, and narrative construction. Within our architecture, decoder-based models are integrated into the RAG pipeline, where they generate context-aware descriptions of attacks by combining classifier outputs with relevant external knowledge.
While proprietary frontier models such as GPT-4 demonstrate remarkable generative capabilities, they cannot be fine-tuned on domain-specific corpora due to the lack of access to their weights. In highly specialized domains such as cybersecurity, this limits their applicability to zero-shot or few-shot prompting, which generally yields lower and less stable results [22, 23]. For this reason, we focus on open-weight encoder models that can be adapted and replicated, while keeping the framework compatible with future open-weight frontier models.
@@ 2.4 AI Agents
The term autonomy refers to the capability and requisites necessary to determine the appropriate course of action to achieve a specified objective. An intelligent agent (AI Agent) is characterized by its ability to perceive its environment, respond to changes within it, initiate actions, and engage with other systems, which may include other agents or human users. One of the core principles foundational to AI agents is the concept of memory. Effective memory management improves an agent’s capability in maintaining contextual awareness. It enables the agent to draw on previous experiences effectively, thereby facilitating the development of incrementally informed decision-making abilities as time progresses. These concepts are depicted in Figure 1. As noted in [25], the emergence of LLMs represents another moment of progress in the realization of AI agents. Substantial advances in this direction have culminated in the emergence of LLM agents. In particular, these agents use LLM as reasoning and planning cores to decide the control flow of an application, while maintaining the characteristics of traditional AI agents. LLM agents extend the capabilities of LLMs by facilitating the use of external tools to address specialized tasks, including mathematical computation and code execution. Ultimately, the model evaluates the adequacy of its output and determines whether additional processing is necessary.
@@ 2.5 Agentic RAG
Although general-purpose LLMs and their associated agents offer remarkable versatility, they frequently lack the in-depth domain-specific expertise essential for addressing intricate and specialized problems. One potential solution to this limitation involves retraining or fine-tuning the model; however, undertaking such procedures can be both financially demanding and resource consuming. An alternative strategy to circumvent these challenges emerges from the RAG paradigm. This strategy enhances LLMs by connecting them to external domain-specific knowledge sources, such as document collections, through a retriever component. Functioning like a search engine, the retriever extracts relevant information from this external repository and integrates it directly into the LLM’s context. This provides the model with the necessary data to effectively answer complex requests.
@@
2.6 Attack classification and detection techniques
In Table 1 some key difference between classical and AI-powered techniques are proposed.
Aspect Classical Techniques AI-powered Techniques (LLMs)
Adaptability to Novel Attacks Low - limited to known patterns or signatures High - capable of generalization and zero-day inference
Example Tools / Frameworks Firewall, IPS/IDS AutoAttacker, SecBench
Work Focus Area Main Contribution Remarks
2.6.1 Classical techniques
In the context of cyber attack classification and detection systems, “classical” techniques generally focus on the definition of taxonomies, attack patterns, and well-established defensive methods. A traditional aspect involves distinguishing between passive attacks (e.g., eavesdropping and traffic analysis) and active attacks (e.g., injecting malicious packets or tampering with data).
An important example of classification is the AVOIDIT taxonomy, which groups attacks according to vectors, objectives, and impacts, underscoring the need for a holistic perspective on security [29]. Within this framework, threats such as Denial of Service (DoS), port scanning, unauthorized data exfiltration, and, more broadly, data integrity violations are analyzed. In parallel, the literature has also focused on distributed attacks (DDoS), where multiple compromised nodes act simultaneously to deny service or cause damage with high impact [34].
On the prevention and response front, many works focus on the integration of cryptographic tools, authentication systems, and key-management protocols to mitigate the risk of device compromise [35]. This approach often assumes that the primary defensive perimeter is defined by cryptographic robustness and the proper configuration of network services, prioritizing the use of IDS and IPS [30, 36].
2.6.2 AI-powered classification and defense techniques
Recent research underscores the growing impact of AI-based tools, particularly Large Language Models , in the domain of cyber-attack orchestration and defense. Early studies demonstrated that LLMs such as ChatGPT could generate attack scripts with success rates ranging from 16% up to 50% when combined with basic cybersecurity skills, thereby lowering the barrier to entry for malicious actors [31, 37]. Offensive platforms like WormGPT and FraudGPT were explicitly developed to harness LLM capabilities for malicious objectives such as phishing, ransomware development, and malware generation [38].
At the same time, the cybersecurity community has begun to leverage LLMs in a defensive capacity, exploiting their ability to interpret contextual nuances in textual data [39, 37]. Advanced frameworks have been proposed to harness LLMs in automated attack classification, attribution, and system hardening. One illustrative example is the AutoAttacker system, which uses LLM-guided modules to orchestrate and evaluate attacks on a simulated organizational network [31].
Studies focusing on threats at the nation-state level also point to the increasing importance of LLM-driven analysis in attributing sophisticated attacks [33]. Furthermore, reviews of cyber defense LLM applications highlight how these models can enhance intrusion detection, automate forensic analysis, and generate real-time alerts, although persistent challenges remain in interpretability, adversarial robustness, and regulatory compliance [38, 37].
A summary of the works discussed above is provided in Table 2.
@@ 3.1 Classification Tool
PD001 1’ and 3580 = ( select count ( * ) from domain.domains as t1, domain.columns as t2, domain.tables as t3 ) -- 0.9999 0.3956 0.0673
PD002 1’’}}{{1016814002+3472965455}} {{’bo’}} {#comment#}{% raw ’bo’.join(’7n’) %} {{’7n’}}{{3140320242+4078937248} 0.3999 0.9997 0.3830
@@ 3.2 RAG Tool
The core LLM engine leverages the semantic representation to automatically generate a natural language query that reflects the identified attack type, salient payload characteristics, and relevant contextual indicators. This query is submitted to a semantic search engine that indexes curated cybersecurity resources, including: (i) vulnerability databases (e.g., Common Vulnerabilities and Exposures CVE, National Vulnerability Database NVD), (ii) technical documentation (e.g., Open Worldwide Application Security Project OWASP, MITRE ATT&CK), (iii) scientific literature and incident reports.
@@ 3.4 User Chat
Although typically hidden, some LLM engines provide access to their internal reasoning processes. In this example, DeepSeek-R1:7B is used, and its reasoning is presented below.